# Analytics, have some humility: a statistical view of fourth-down decision making

### v5

@@ 4.1 Example plays
Example play 1. First, recall example play 1 from Section 2.2 in which the Patriots had the ball against the Colts in Week 10 of 2023. In Figure 5 we knit uncertainty quantification into decision making. Although the point estimate (the blue column in Figure 5a) suggests that ${\mathsf{Go}}$ provides a $1.5\%$ gain in win probability over ${\mathsf{FG}}$, $40\%$ of the bootstrapped models view ${\mathsf{FG}}$ as better than ${\mathsf{Go}}$ (the orange column in Figure 5a), reflecting substantial uncertainty in the optimal fourth-down decision. Also, the $90\%$ confidence interval is $[-6\%,6\%]$, and the histogram of bootstrapped win probability gain estimates in Figure 5b is centered near zero and has a large spread. These suggest that ${\mathsf{Go}}$ could either be a good or a bad decision.
@@
Figure 6: Decision boundary charts for example play 2. Figure (a) visualizes the estimated optimal decision (color) according to effect size (color intensity, where darker colors indicate larger values) as a function of yardline ($x$-axis) and yards to go ($y$-axis), holding the other game-state variables constant. The pink dot represents the actual play’s yardline and yards to go. Figure (b) is similar except color intensity reflects bootstrap percentage.
We include more examples of this fourth-down decision procedure in Appendix E. We also include an interactive Shiny App that visualizes the decision procedure for any game-state on Github.555 To run the Shiny app, download the Github repository at https://github.com/snoopryan123/fourth_down and run 3_shiny/app.R.
@@ 4.2 Quantifying overconfidence in win probability point estimates
In Figure 7 we visualize the distributions of effect size and uncertainty across all fourth-down plays from 2018-2022. Effect size is the estimated gain in win probability and uncertainty is proportional to bootstrap percentage. To facilitate easier communication to a non-technical football coach, we bin decision confidence levels into three buckets: confident (${\mathsf{boot\%}}\in[90\%,100\%]$), lean (${\mathsf{boot\%}}\in[75\%,90\%)$), and uncertain (${\mathsf{boot\%}}\in[0\%,75\%]$).
@@
We are confident in the vast majority of decisions that have an effect size above $4\%$. Most decisions with an effect size under $1\%$ are uncertain, and decisions with an effect size between $1\%$ and $4\%$ are a healthy split between confident, lean, and uncertain. Many plays ($45\%$) have an effect size under $1\%$ and most plays ($87\%$) have an effect size under $4\%$. So, most fourth-down plays feature an effect size that is subject to considerable uncertainty. In particular, we are confident in just $30\%$ of all fourth-down decisions from 2018 to 2022 and $44\%$ of them are uncertain. This analysis reflects substantial overconfidence in win probability point estimates; far fewer fourth-down decisions are as obvious as traditional analyses suggest.
Interestingly, we are confident in about $10\%$ of decisions with effect size under $1\%$ and in about $20\%$ of decisions with effect size between $1\%$ and $2\%$. Although decision confidence is correlated with estimated gain in win probability, they are fundamentally different: we can be confident in plays that provide small edges. Over the course of a season, these edges can accumulate into a large advantage that coaches should take advantage of.
Furthermore, a football analyst shouldn’t penalize a coach for making a decision that has high uncertainty regardless of the effect size because the estimated edge may be due to noise. Accordingly, we should evaluate coaches only on plays for which we are confident (and perhaps also plays in which we lean towards a decision). In Figure 8 we rank coaches by the proportion of fourth-down decisions they made in accordance with our model from 2018 to 2022 among fourth-down plays we are confident in. The top of this list has coaches like John Harbaugh from “analytics bent” organizations but also coaches who are considered more traditional on fourth down such as Ron Rivera, Sean McVay, and Mike Vrabel. While “analytics bent” coaches tend to follow fourth-down recommendations from win probability point estimates, which include recommendations on confident plays, some of the more traditional coaches seem to have a good feel for obvious fourth downs.
@@
Despite overconfidence in win probability point estimates, analysts have been largely correct that NFL coaches do not go for it enough on fourth down. Across all fourth-down plays from 2018-2022 that we are confident in, the coach made the right decision for $90\%$ of plays where they should have kicked but just $57\%$ of plays where they should have gone for it. Play calling in the NFL is still far too conservative: coaches consistently make wrong decisions, particularly when they should go for it.
@@ 5 Discussion
Win probability models are commonly used across sports analytics. Fourth-down decision making is a prominent example. There has been little reflection on the high-variance nature of win probability estimators and no substantial attempt to quantify uncertainty in these models. In this study, we argue it is imperative to knit uncertainty quantification into win probability-based decision making. Using fourth-downs as our case study, we in particular recommend a fourth-down decision when we are confident it has higher win probability than all other decisions. If we are not confident in a recommendation because it is subject to substantial uncertainty, we should not penalize a coach who doesn’t follow it. Similarly, we do not think it is right for a football analyst to recommend an uncertain fourth-down recommendation, regardless of the effect size. Despite analysts’ overconfidence in win probability point estimates, after accounting for uncertainty we still find that NFL coaches are too conservative: they don’t go for it enough on fourth down.
Our analysis is not without limitations and there are many avenues for future work. To begin, in this study, we used bootstrapping to quantify uncertainty in win probability estimates. Although our bootstrap procedure produces wide confidence intervals, it underestimates uncertainty since it quantifies sampling uncertainty but not model uncertainty. The former is uncertainty resulting from fitting a model on a finite dataset (“variance”) and the latter is uncertainty caused by our model being wrong or biased (“bias”). Our models are subject to several sources of model uncertainty. We made several simplyfing assumptions in modeling fourth-down win probability, such as swapping the expectation with first-down win probability. Furthermore, “true” win probability is a function of unobserved confounders. For example, how well Tom Brady slept the previous night could affect his team’s win probability. A more elaborate analysis would capture model uncertainty.
@@ Appendix A Data details
posteam spread pre-game point spread relative to the possession team
yardline num. yards from the opponent’s endzone in $\{0,1,...,99,100\}$ (00 is touchdown and $100$ is safety)
ydstogo the number of yards the offense has left to gain before achieving a first down or a touchdown
@@ Appendix B Estimating player / team quality
Punter quality. We define a punter’s quality by a weighted sum of his punt yards over expected over all his previous punts in his career. To begin, we fit a simple expected next yardline after punting model ${{\mathbb{E}}_{{\mathsf{Punt}}}^{(0)}}$ on a held-out dataset of all punts from 1999 to 2005. We fit ${{\mathbb{E}}_{{\mathsf{Punt}}}^{(0)}}$ using linear regression as a function of yardline (specifically, a cubic polynomial in yardline). Then, we define the punt yards over expected (${\mathsf{PYOE}}$) of the $n^{th}$ punt by
@@
Now we define punter quality using Rams’ punter Johnny Hekker for concreteness. Index all of Hekker’s punts from 2006 to 2021 by $n$. We define Hekker’s punter quality prior to punt $n$ by a weighted sum of the punt yards over expected from his previous kicks,
${\mathsf{pq}}_{n}:=\alpha\cdot{\mathsf{pq}}_{n-1}+{\mathsf{PYOE}}_{n-1},$ (22)
where ${\mathsf{pq}}_{0}:=0$ and ${\mathsf{PYOE}}_{0}:=0$. We use an exponential decay weight $\alpha=0.995$ to upweight more recent punts. For instance, a play which occurred 138 plays ago is weighted half as much as the previous play since $0.50.995^{138}=$ 0.5. Also, a play which occurred 459 plays ago is weighted one-tenth as much as the previous play since $0.10.995^{459}=$ 0.1. Finally, we standardize the punter quality column to have mean zero and standard deviation $1/2$. In Figure 9a we plot the career mean punter quality of each punter with over 250 punt attempts from 2006 to 2021. As expected, Johnny Hekker has the highest punter quality.
Figure 9: (a) The career mean punter quality ($x$-axis) of each punter ($y$-axis) with over 250 punts from 2006 to 2021. (b) The career mean kicker quality ($x$-axis) of each kicker ($y$-axis) with over 100 field goal attempts from 2006 to 2021.
Kicker quality. We define a kicker’s quality by a weighted sum of his field goal probability added over all his previous kicks in his career. To begin, we fit a simple field goal probability model ${P_{\mathsf{FG}}^{(0)}}$ on a held-out dataset of all field goals from 1999 to 2005. We fit ${P_{\mathsf{FG}}^{(0)}}$ using logistic regression as a function of yardline (specifically, a cubic spline with five degrees of freedom on the yardline). Then, we define the field goal probability added (${\mathsf{FGPA}}$) of the $n^{th}$ field goal by
Now we define kicker quality using the Ravens’ kicker Justin Tucker for concreteness. Index all of Tucker’s field goals from 2006 to 2021 by $n$. We define Tucker’s kicker quality prior to field goal $n$ by a weighted sum of the field goal probability added in his previous kicks,
${\mathsf{kq}}_{n}:=\alpha\cdot{\mathsf{kq}}_{n-1}+{\mathsf{FGPA}}_{n-1},$ (24)
where ${\mathsf{kq}}_{0}:=0$ and ${\mathsf{FGPA}}_{0}:=0$. As before, we use an exponential decay weight $\alpha=0.995$ to upweight more recent kicks. Finally, we standardize the kicker quality column to have mean zero and standard deviation $1/2$. In Figure 9b we plot the career mean kicker quality of each kicker with over 100 field goal attempts from 2006 to 2021. As expected, Justin Tucker has by far the highest kicker quality.
Team quality. We create measures of offensive and defensive quality using the result of each previous play. A common and good way of quantifying play success is expected points added (${\mathsf{EPA}}$) (Yurko et al.,, 2018). The ${\mathsf{EPA}}$ of play $n$ is the difference in expected points (${\mathsf{EP}}$) between the end and the beginning of the play,
${\mathsf{EPA}}_{n}={\mathsf{EP}}_{n}-{\mathsf{EP}}_{n-1}.$ (25)
We remove all plays from 1999 to 2005 from our play-by-play dataset and fit an expected points model ${\mathsf{EP^{(0)}}}$ to this held-out data. In fitting this model, the outcome variable is the observed net points of the next score, a real number in $\{7,3,2,0,-2,-3,-7\}$. We model expected net points as a linear function of yardline, log yards to go, and one combined indicator for third and fourth down.
Figure 10: The career mean quarterback quality ($x$-axis) of each quarterback ($y$-axis) with over 1750 attempts and whose rookie season came after 2006.
Then we use ${\mathsf{EPA^{(0)}}}$ (${\mathsf{EPA}}$ derived from ${\mathsf{EP^{(0)}}}$) to craft our team and player quality metrics. For concreteness, consider deriving the quarterback quality of Patrick Mahomes. Index all the plays from 2006 to 2021 in which Mahomes passes or runs the ball by $n$. We define Mahomes’ quarterback quality prior to play $n$ by a weighted sum of the ${\mathsf{EPA^{(0)}}}$ of his previous plays,
$q_{n}:=\gamma_{n}\cdot\alpha\cdot q_{n-1}+{\mathsf{EPA^{(0)}}}_{n-1},$ (26)
where $q_{0}:=0$ and ${\mathsf{EPA^{(0)}}}_{0}:=0$. We use an exponential decay weight $\alpha$ to upweight more recent plays. We set $\alpha=0.995$ for each of our team and player quality metrics to upweight more recent plays. Additionally, at the start of each season, we shrink by a multiplicative factor $\gamma$. So, the shrinkage weight $\gamma_{n}$ of play $n$ is $\gamma$ if this is Mahomes’ first play of the season, otherwise it is 1. We use different shrinkage values $\gamma$ for each team and player quality metric, shown in Table 3. We tuned these $\gamma$ values to optimize predictive performance on a held-out validation set. A smaller $\gamma$ reflects a quantity that is noisier from year to year. Finally, we standardize each offensive and defensive quality metric to have mean zero and standard deviation $1/2$. In Figure 10 we plot the career mean quarterback quality of each quarterback with over 1750 attempts and whose rookie season came after 2006. As expected, Patrick Mahomes has by far the highest quarterback quality. We construct the other team quality metrics described in Table 3 in a similar fashion, via Formula (26).
variable $\gamma$
quarterback quality $3/4$
non-quarterback offensive quality $1/2$
defensive quality against the pass $1/3$
defensive quality against the run $1/3$
Table 3: Summary of team quality metrics computed via Formula (26) and their start-of-season shrinkage parameters $\gamma$.

### v6

@@ 4.1 Example plays
Example play 1. First, recall example play 1 from Section 2.2 in which the Patriots had the ball against the Colts in Week 10 of 2023. In Figure 5 we knit uncertainty quantification into decision making. Although the point estimate (the blue column in Figure 5a) suggests that ${\mathsf{Go}}$ provides a $1.1\%$ gain in win probability over ${\mathsf{FG}}$, $44\%$ of the bootstrapped models view ${\mathsf{FG}}$ as better than ${\mathsf{Go}}$ (the orange column in Figure 5a), reflecting substantial uncertainty in the optimal fourth-down decision. Also, the $90\%$ confidence interval is $[-4\%,5\%]$, and the histogram of bootstrapped win probability gain estimates in Figure 5b is centered near zero and has a large spread. These suggest that ${\mathsf{Go}}$ could either be a good or a bad decision.
@@
Figure 6: Decision boundary charts for example play 2. Figure (a) visualizes the estimated optimal decision (color) according to effect size (color intensity, where darker colors indicate larger values) as a function of yards to opponent endzone ($x$-axis) and yards to go ($y$-axis), holding the other game-state variables constant. The pink dot represents the actual play’s yards to opponent endzone and yards to go. Figure (b) is similar except color intensity reflects bootstrap percentage.
We include more examples of this fourth-down decision procedure in Appendix E. We also include an interactive Shiny App that visualizes the decision procedure for any game-state on Github.666 To run the Shiny app, download the Github repository at https://github.com/snoopryan123/fourth_down and run 3_shiny/app.R.
@@ 4.2 Quantifying overconfidence in win probability point estimates
In Figure 7 we visualize the distributions of effect size and uncertainty across all fourth-down plays from 2018-2022. Effect size is the estimated gain in win probability and uncertainty is proportional to bootstrap percentage. To facilitate easier communication to a non-technical football coach, we bin decision confidence levels. We partition $[50\%,100\%]$ into three equally spaced buckets and add $[0\%,50\%)$ to the lowest bucket: confident (${\mathsf{boot\%}}\in[83\%,100\%]$), lean (${\mathsf{boot\%}}\in[67\%,83\%)$), and uncertain (${\mathsf{boot\%}}\in[0\%,67\%)$).
@@
We are confident in the vast majority of decisions that have an effect size above $4\%$. Most decisions with an effect size under $1\%$ are uncertain, and decisions with an effect size between $1\%$ and $4\%$ are a healthy split between confident, lean, and uncertain. Many plays ($44\%$) have an effect size under $1\%$ and most plays ($87\%$) have an effect size under $4\%$. So, many fourth-down plays feature an effect size that is subject to considerable uncertainty. In particular, we are confident in just $48\%$ of all fourth-down decisions from 2018 to 2022 and $27\%$ of them are uncertain. This analysis reflects substantial overconfidence in win probability point estimates; far fewer fourth-down decisions are as obvious as traditional analyses suggest.
Interestingly, we are confident in about $20\%$ of decisions with effect size under $1\%$ and in about $50\%$ of decisions with effect size between $1\%$ and $2\%$. Although decision confidence is correlated with estimated gain in win probability, they are fundamentally different: we can be confident in plays that provide small edges. Over the course of a season, these edges can accumulate into a large advantage that coaches should take advantage of.
Furthermore, a football analyst shouldn’t penalize a coach for making a decision that has high uncertainty regardless of the effect size because the estimated edge may be due to noise. Accordingly, we should evaluate coaches only on plays for which we are confident (and perhaps also plays in which we lean towards a decision). In Figure 8 we rank coaches by the proportion of fourth-down decisions they made in accordance with our model from 2018 to 2022 among fourth-down plays we are confident in. The top of this list has coaches like John Harbaugh from “analytics bent” organizations but also coaches who are considered more traditional on fourth down such as Ron Rivera and Pat Shurmur. While “analytics bent” coaches tend to follow fourth-down recommendations from win probability point estimates, which include recommendations on confident plays, some of the more traditional coaches seem to have a good feel for obvious fourth downs.
@@
Despite overconfidence in win probability point estimates, analysts have been largely correct that NFL coaches do not go for it enough on fourth down. Across all fourth-down plays from 2018-2022 that we are confident in, the coach made the right decision for $91\%$ of plays where they should have kicked but just $49\%$ of plays where they should have gone for it. Play calling in the NFL is still far too conservative: coaches consistently make wrong decisions, particularly when they should go for it.
@@ 5 Discussion
Win probability models are commonly used across sports analytics. Fourth-down decision making is a prominent example. There has been little reflection on the high-variance nature of win probability estimators and no substantial attempt to quantify uncertainty in these models. In this study, we argue it is imperative to knit uncertainty quantification into win probability-based decision making. Using fourth-downs as our case study, we in particular recommend a fourth-down decision when we are confident it has higher win probability than all other decisions. If we are not confident in a recommendation because it is subject to substantial uncertainty, we should not penalize a coach who doesn’t follow it. Similarly, we do not think it is right for a football analyst to recommend an uncertain fourth-down recommendation, regardless of the effect size. Despite analysts’ overconfidence in win probability point estimates, after accounting for uncertainty we still find that NFL coaches are too conservative: they don’t go for it enough on fourth down. Additionally, as teams across the NFL have increased their propensity to go for it on fourth down, the success rate hasn’t declined.777 https://medium.com/@dacr444/analyzing-4th-down-attempts-over-the-last-24-seasons-8c0f99c538cb).
Our analysis is not without limitations and there are many avenues for future work. To begin, in this study, we used bootstrapping to quantify uncertainty in win probability estimates. Although our bootstrap procedure produces wide confidence intervals, it underestimates uncertainty since it quantifies sampling uncertainty but not model uncertainty. The former is uncertainty resulting from fitting a model on a finite dataset (“variance”) and the latter is uncertainty caused by our model being wrong or biased (“bias”). Our models are subject to several sources of model uncertainty. We made several simplyfing assumptions in modeling fourth-down win probability, such as swapping the expectation with first-down win probability. Furthermore, “true” win probability is a function of unobserved confounders. For example, the yards to go variable would ideally be derived from tracking data and measured in inches, but the publicly available version is integer-valued. Lopez, (2020) shows that, particularly at short distances (i.e., comparing $4^{th}$ down and inches to $4^{th}$ and 1), a more fine-grained yards to go variable can produce substantially different forecasts. Another key unobserved confounder is the play call that’s being considered for a potential conversion attempt. For instance, on $4^{th}$ and 1, the Eagles “tush push” play call has a much higher success rate than the league-wide base rate.888 https://www.espn.com/nfl/story/_/id/41234531/philadelphia-eagles-tush-push-jason-kelce The play call can make or break a conversion attempt, and not adjusting for it introduces additional uncertainty into the analysis. Unfortunately, it is not publicly available. A more elaborate analysis would capture model uncertainty.
@@ Appendix A Data details
posteam spread pre-game Vegas point spread relative to the possession team
total points line pre-game Vegas total points over/under line
yards to opp. endzone (i.e., yardline) num. yards to the opponent’s endzone in $\{0,1,...,99,100\}$ (00 is touchdown and $100$ is safety)
yards to go (i.e., ydstogo) an integer, the number of yards an offense has to gain to achieve a first down or touchdown
@@ Appendix B Estimating player / team quality
Kicker quality. We define a kicker’s quality by a weighted mean of his field goal probability added over all his previous kicks in his career. To begin, we fit a simple kicker-agnostic field goal probability model ${P_{\mathsf{FG}}^{(0)}}$ using logistic regression as a function of yardline (specifically, a cubic polynomial in yardline). Then, we define the field goal probability added (${\mathsf{FGPA}}$) of the $n^{th}$ field goal by
Now we define kicker quality using the Ravens’ kicker Justin Tucker for concreteness. Index all of Tucker’s field goals by $n$. We define Tucker’s kicker quality prior to field goal $n\geq 2$ by a weighted mean of the field goal probability added in his previous kicks,
${\mathsf{kq}}_{n}:=\frac{\sum_{j=1}^{n-1}\alpha^{n-1-j}\cdot{\mathsf{FGPA}}_{j}}{\gamma+\sum_{j=1}^{n-1}\alpha^{n-1-j}},$ (24)
and ${\mathsf{kq}}_{1}:=0$. The hyperparameter $\gamma$ stabilizes ${\mathsf{kq}}$ for low $n$, shrinking ${\mathsf{kq}}$ towards 0 (i.e., shrinking towards an average kicker). In other words, we impute $\gamma$ synthetic field goals, each with an “average” ${\mathsf{FGPA}}$ outcome of zero. We use $\gamma=96$. The hyperparameter $\alpha$ is an exponential decay weight, which upweights more recent kicks, allowing us to capture non-stationarity in kicker quality over time. We use $\alpha=0.985$. For instance, a kick which occurred 46 kicks ago is weighted half as much as the previous kick since $0.985^{46-1}\approx 0.5$. Also, a kick which occurred 153 kicks ago is weighted one-tenth as much as the previous kick since $0.985^{153-1}\approx 0.1$. Finally, we standardize kicker quality to have mean zero and standard deviation $1$.
In Figure 9 we visualize the kicker quality trajectories of a few kickers across their careers. We see non-stationarity in kicker quality over time. For example, notice Younghoe Koo’s sharp rise in 2020 (he made the Pro Bowl that year) and Robbie Gould’s dip in 2019. In Figure 10a we plot the career mean kicker quality of each kicker with over 100 field goal attempts in our dataset. As expected, Justin Tucker has by far the highest kicker quality.
Figure 9: Kicker quality trajectories versus time for various kickers.
Figure 10: (a) The career mean kicker quality ($x$-axis) of each kicker ($y$-axis) with over 100 field goal attempts in our dataset. (b) The career mean punter quality ($x$-axis) of each punter ($y$-axis) with over 250 punts in our dataset.
Punter quality. We define punter quality in a similar fashion as kicker quality. In particular, we define a punter’s quality by a weighted mean of his punt yards over expected over all his previous punts in his career. To begin, we fit a simple punter-agnostic expected next yardline after punting model ${{\mathbb{E}}_{{\mathsf{Punt}}}^{(0)}}$ using linear regression as a function of yardline (specifically, a cubic polynomial in yardline). Then, we define the punt yards over expected (${\mathsf{PYOE}}$) of the $n^{th}$ punt by
@@
Now we define punter quality using Rams’ punter Johnny Hekker for concreteness. Index all of Hekker’s punts by $n$. We define Hekker’s punter quality prior to punt $n\geq 2$ by a weighted mean of the punt yards over expected in his previous kicks,
${\mathsf{pq}}_{n}:=\frac{\sum_{j=1}^{n-1}\alpha^{n-1-j}\cdot{\mathsf{PYOE}}_{j}}{\gamma+\sum_{j=1}^{n-1}\alpha^{n-1-j}},$ (26)
and ${\mathsf{pq}}_{1}:=0$. The hyperparameters $\gamma$ and $\alpha$ play the same role as in kicker quality. We use $\gamma=150$ and $\alpha=0.99$. Finally, we standardize punter quality to have mean zero and standard deviation $1$.
In Figure 11 we visualize the punter quality trajectories of a few punters across their careers. We see that non-stationarity in punter quality over time. For example, notice Pat McAfee’s steady rise across his mid and later years (he made the Pro Bowl in 2014 and 2016) and Jake Bailey’s decline in 2022. In Figure 10b we plot the career mean punter quality of each punter with over 250 punts in our dataset. As expected, Johnny Hekker has the highest punter quality.
Figure 11: Punter quality trajectories versus time for various punters.
Difference in offensive and defensive quality. We create a measure of how much better the offensive team’s offensive quality is than the defensive team’s defensive quality using information from betting markets. We use Vegas’ pre-game point spread relative to the offensive team (denoted ${\mathsf{PS}}$) and the total points over/under line (denoted ${\mathsf{TP}}$). In particular, our measure is $({\mathsf{TP}}-{\mathsf{PS}})/2$, standardized to have mean zero and standard deviation $1$.
We derive this expression as follows. Denote the offensive quality of the offensive team by ${\mathsf{oqot}}$, the defensive quality of the defensive team by ${\mathsf{dqdt}}$, the offensive quality of the defensive team by ${\mathsf{oqdt}}$, and the defensive quality of the offensive team by ${\mathsf{dqot}}$. These variables are not observable, and we are interested in the differences ${\mathsf{oqot}}-{\mathsf{dqdt}}$ and ${\mathsf{oqdt}}-{\mathsf{dqot}}$. The point spread is a measure of how much better the offensive team’s offense’s edge is over the opposing defense relative to the opposing team’s edge. A reasonable model is
${\mathsf{PS}}=-(({\mathsf{oqot}}-{\mathsf{dqdt}})-({\mathsf{oqdt}}-{\mathsf{dqot}})),$ (27)
recalling that, by convention, negative point spreads indicate larger edges. The total points line is a measure of the combined strenth of both offense’s edges. A reasonable model is
${\mathsf{TP}}=({\mathsf{oqot}}-{\mathsf{dqdt}})+({\mathsf{oqdt}}-{\mathsf{dqot}})+\overline{{\mathsf{TP}}},$ (28)
where $\overline{{\mathsf{TP}}}$ denotes the average total points line if neither offense has an edge. From these two equations, we derive ${\mathsf{oqot}}-{\mathsf{dqdt}}=({\mathsf{TP}}-{\mathsf{PS}})/2$ and ${\mathsf{oqdt}}-{\mathsf{dqot}}=({\mathsf{TP}}+{\mathsf{PS}})/2$, ignoring an additive constant.